# SENSE models: an open source solution for multilingual and multimodal semantic-based tasks

### v1

@@ VI-B2 Slot Filling
Table VII reports the COER and CVER on three SLU datasets using full fine-tuning of four models. Across all settings, SENSE consistently yields the lowest error rates, outperforming w2v-BERT 2.0, SAMU-XLSR, and SONAR.
TABLE VII: COER and CVER (%) for slot filling task (FR, EN, and TN)
@@ VII Frame‐Level Semantic Attention Analysis
Figure 3 shows the curve of attention weights applied to each frame embedding computed by the SENSE model for a single utterance. We can see the high peak at the beginning, as well as additional peaks appearing in regions where Figure 2 was relatively flat. Additionally, the VAD curve (in orange) indicates the presence of non‐speech segments. We continue our investigation by applying a voice activity detector (VAD) to the audio files. For this purpose, we relied on the Silero VAD tool [43]. We also applied WhisperX [44] to obtain forced word alignments. Although these word alignments are not very precise, they provide useful information for understanding how semantics are captured locally.
@@
The visualization of the word alignment helps explain the peaks observed in the attention curve: they usually correspond to the beginning of content words with high semantic information, while function words (determiners, some prepositions, etc.) do not trigger such behavior.
To further investigate multilingual aspects beyond French, we also examine the German utterances derived from the French ones in the VoxPopuli corpus. We observe similar behavior of the SENSE model across both languages, as illustrated in Figure 4, with peaks of attention at the beginning of the utterances, and at the beginning of some real words. The utterance in Figure 4 is the German spoken translation of the French utterance in Figure 3.

### v2

@@ VI-B2 Slot Filling
Table VII reports the COER and CVER on four SLU datasets using full fine-tuning of four models. Across all settings, SENSE consistently yields the lowest error rates, outperforming w2v-BERT 2.0, SAMU-XLSR, and SONAR.
TABLE VII: COER and CVER (%) for slot filling task
@@ VII Frame‐Level Semantic Attention Analysis
Figure 3 shows the curve of attention weights applied to each frame embedding computed by the SENSE model for a single utterance. We can see the high peak at the beginning, as well as additional peaks appearing in regions where Figure 2 was relatively flat. Additionally, the VAD curve (in orange) indicates the presence of non‐speech segments. We continue our investigation by applying a voice activity detector (VAD) to the audio files. For this purpose, we relied on the Silero VAD tool [SileroVAD]. We also applied WhisperX [bain2022whisperx] to obtain forced word alignments. Although these word alignments are not very precise, they provide useful information for understanding how semantics are captured locally.
@@
The visualization of the word alignment helps explain the peaks observed in the attention curve: they usually correspond to the beginning of content words with high semantic information, while function words (determiners, some prepositions, etc.) do not trigger such behavior. To further investigate multilingual aspects beyond French, we also examine the German utterances derived from the French ones in the VoxPopuli corpus. We observe similar behavior of the SENSE model across both languages, as illustrated in Figure 4, with peaks of attention at the beginning of the utterances, and at the beginning of some real words. The utterance in Figure 4 is the German spoken translation of the French utterance in Figure 3.
@@ Acknowledgment
-A List of languages and duration used to train SENSE
This section presents the list of languages used to train the SENSE model, along with their corresponding ISO codes and total durations of available speech data. The dataset covers a diverse range of languages with varying amounts of audio, as summarized in Table X.
TABLE X: Training Data Duration by Language
Lang ISO Train Lang ISO Train
af afr 00:11:08 am amh 00:48:24
ar arb 32:21:50 as asm 01:05:47
ast 00:30:18 ba bashk 138:28:27
be bel 472:35:42 bg bul 07:05:32
bn ben 32:46:34 br bre 02:28:24
ca cat 1748:12:47 ckb 08:59:10
cs ces 27:13:10 cv chv 01:59:10
cy cym 11:26:32 da dan 04:07:44
de deu 928:06:24 dv div 03:46:30
el ell 02:07:56 en eng 1743:18:52
eo epo 235:04:29 es spa 496:25:19
et est 05:42:40 fa fas 29:53:59
fi fin 02:18:01 fr fra 814:54:55
fy-NL fry 05:31:05 ga-IE gle 00:35:53
gl glg 42:00:53 gn grn 01:20:21
he heb 01:22:10 hi hin 05:47:27
hsb 01:27:17 hu hun 53:41:56
ia ina 04:30:53 id ind 07:38:05
it ita 251:08:54 ja jpn 13:03:50
ka kat 86:44:37 kab 143:29:23
$k$ kaz 00:44:12 ko kor 00:50:02
ky kir 02:18:33 lt lit 10:06:25
lv lav 21:14:15 ml mal 01:21:47
mn khk 03:06:08 mhr 209:15:15
mk mkd 02:05:56 mr mar 01:49:35
mt mlt 02:23:04 ne-NP npi 00:21:48
nl nld 45:29:51 $n$-NO nno 00:35:44
oc oci 00:22:34 or ory 02:40:10
os oss 00:17:07 pa-IN pan 01:09:32
pl pol 29:47:25 ps pus 02:23:12
pt por 24:35:58 ro ron 05:40:03
ru rus 37:35:58 sah 03:04:22
sc srd 00:44:20 sk slk 03:31:20
sl slv 01:20:46 sr srp 00:36:42
sv-SE swe 08:47:01 sw swh 69:18:50
ta tam 68:08:48 th tha 37:09:31
ti tir 00:01:16 tk tuk 01:03:33
tr tur 40:04:47 $t$ tat 09:14:22
ug uig 80:12:36 uk ukr 30:22:51
ur urd 08:07:29 uz uzn 52:31:28
vi vie 02:49:31 yo yor 02:05:47
zh-HK zho 09:29:54
-B Details of hyperparams for downstream tasks
-B1 Slot Filling and Named Entity Recognition
For the fine-tuning phase, we adopt as a downstream probe a straightforward architecture consisting of a one layer DNN followed by a softmax activation function. The training is performed using the Connectionist Temporal Classification (CTC) loss. The probe’s hidden layers have a dimension of 1024, with a dropout rate of 0.15. We use different learning rates and optimizers for the pre-trained encoder and the probe. The speech encoder is fine-tuned using a learning rate of $1\times 10^{-5}$ with a probe learning rate of $1.5$. The batch size is 4. We fine-tune the entire models for 50 epochs.
-B2 Speech summarization
For the task of speech summarization, we trained the BARThez model to generate summaries from speech embeddings. However, the BARThez model expects input sequences as vectors of size 768, whereas the embedding models we integrate may produce vectors of different sizes. To align these representations with BARThez’s expected input space, we introduce a linear projection layer followed by a GeLU activation function, except for SAMU-XLSR embeddings, which already match BARThez’s dimensionality and thus require only the activation function. The first training stage uses the French MLSUM dataset, as it is one of the biggest summarization datasets available in French. This step is essential for helping the model adapt to sentence embeddings as input. We trained our model using the AdamW optimizer, with a batchsize of 16 and different learning rates for the pretrained seq2seq model and the projection layer. The pretrained seq2seq model parameters were optimized with a learning rate of $1\times 10^{-5}$, while the linear projection layer was trained with a larger learning rate of $1\times 10^{-3}$, allowing faster adaptation of the randomly initialized weights. A weight decay of $1\times 10^{-5}$ was applied uniformly to regularize both components. We used the standard cross-entropy loss to train the model on this sequence prediction task.
The second training stage specializes the model on task-specific dataset. Thus, for speech summarization, we continue the training of the model on DECODA and evaluated the performance by using speech utterance embeddings generated from the test-set. As the goal is to compare SONAR, SAMU-XLSR and SENSE embeddings in this task, we did the following trainings :
  • We trained the BARThez model on the MLSUM dataset using text sentence embeddings generated by the SONAR text encoder, and subsequently fine-tuned it on the DECODA dataset using SONAR speech embeddings.
  • We trained the BARThez model on the MLSUM dataset using text sentence embeddings from the LaBSE text encoder, and then fine-tuned it on DECODA using SAMU-XLSR speech embeddings.
  • We trained the BARThez model on the MLSUM dataset using text sentence embeddings from the BGE text encoder, followed by fine-tuning on DECODA using SENSE speech embeddings.
-B3 Speech Translation
We initialize the encoder with a speech encoder and decoder using mBart [liu2020mbart]. A feed-forward network projection layer is used to connect the encoder and decoder, bridging the two modules. The speech encoder is fine-tuned with a learning rate of $1\times 10^{-4}$, and the mBART decoder with $1\times 10^{-4}$, while the feed-forward projection layer before the decoder is trained with a higher learning rate of $1\times 10^{-3}$. All modules are optimized using the Adam optimizer. We train the model over 100 epochs with a batch size of 4.